### v1

# SG-LRA: Self-Generating Automatic Scoliosis Cobb Angle Measurement with Low-Rank Approximation

Abstract
Automatic Cobb angle measurement from X-ray images is crucial for scoliosis screening and diagnosis. However, most existing regression-based methods and segmentation-based methods struggle with inaccurate spine representations or mask connectivity/fragmentation issues. Besides, landmark-based methods suffer from insufficient training data and annotations. To address these challenges, we propose a novel framework including Self-Generation pipeline and Low-Rank Approximation representation (SG-LRA) for automatic Cobb angle measurement. Specifically, we propose a parameterized spine contour representation based on LRA, which enables eigen-spine decomposition and spine contour reconstruction. We can directly obtain spine contour with only regressed LRA coefficients, which form a more accurate spine representation than rectangular boxes. Also, we combine LRA coefficient regression with anchor box classification to solve inaccurate predictions and mask connectivity issues. Moreover, we develop a data engine with automatic annotation and automatic selection in an iterative manner, which is trained on a private Spinal2023 dataset. With our data engine, we generate the largest scoliosis X-ray dataset named Spinal-AI2024 largely without privacy leaks. Extensive experiments on public AASCE2019, private Spinal2023, and generated Spinal-AI2024 datasets demonstrate that our method achieves state-of-the-art Cobb angle measurement performance. Our code and Spinal-AI2024 dataset are available at https://github.com/Ernestchenchen/SG-LRA and https://github.com/Ernestchenchen/Spinal-AI2024, respectively.
{IEEEkeywords}
Automatic Cobb angle measurement, self-generation pipeline, low-rank approximation representation, largest scoliosis X-ray dataset
1 Introduction
[…]
Scoliosis is an important spinal disease in human beings, especially for Scoliosis, a common spinal deformity that primarily affects adolescents, can significantly impact the health of patients [1, 2, 3]. Early detection and treatment of scoliosis can help reduce the need for surgical interventions [4]. The Cobb angle [5] is one of the most widely used clinical measures for evaluating and diagnosing the severity of scoliosis. It is defined as the angle between parallel lines of the top of the upper end vertebra (the most tilted vertebrae on the curve) and the bottom of the lower one [6]. There are three regional spinal deformities associated with different spinal segments: proximal thoracic (PT), main thoracic (MT), and thoracolumbar/lumbar (TL/L), as illustrated in Fig. 1(c).
Figure 1: Illustration of Cobb angle measurement. There are three categories of scoliosis: proximal thoracic (PT) curve, main thoracic (MT) curve, and thoracolumbar/lumbar (TL/L) curve, as shown in (c). In (a), handwritten 23°, 31°, and 34° are manually measured by experts, angles in green are measured based on rectangular boxes (also in green), and angles in red are measured based on our proposed contour boxes composed by spinal landmarks (drawn in (b)). (b) and (c) show two types of evaluations: maximum angle of three spinal curves, three regional angles, respectively.
Currently, manual Cobb angle measurement on radiographic images like X-ray images by experts is still the mainstream way in medical institutions. However, it is time-consuming and often has potential errors [7, 8], since it can be affected by spinal selection and inter-observer variability, which can be observed from the distances between manual angles and angles by landmarks in Fig. 1(a). It has been validated that manual annotation of the Cobb angle exhibits fluctuations ranging from 2.8 to 8 degrees [9]. Therefore, developing a reliable automatic Cobb angle measurement method is crucial.
Traditional automatic Cobb angle measurement methods rely on hand-crafted features [10, 11], limiting model performance and generalization. In recent years, some regression-based methods use rectangular boxes as the spine representation [12, 13], which suffer from reduced accuracy when spinal blocks are wedge-shaped or irregular, as shown in Fig. 1(a). Segmentation-based methods are sensitive to image quality [14, 15], tending to predict connected or damaged spinal masks for low-quality inputs. Landmark-based methods necessitate vast amounts of training data and annotations [16, 17, 18], which are costly and time-consuming to collect and label, and are exacerbated by data scarcity due to patient privacy concerns.
In this paper, we distill the above limitations as three main issues, and tackle them from three aspects: (a) How to achieve more accurate spine representation? (b) How to generate sufficient scoliosis images while protecting patient privacy? (c) How to obtain accurate annotations at low costs?
Parameterized Spine Representation. We propose a novel automatic Cobb angle measurement framework named SG-LRA. We combine regression-based and landmark-based techniques, by using a set of spinal landmarks to represent the bounding box of spinal segments. Inspired by low-rank approximation (LRA) parameterized contour representation in object detection [19, 20], we use LRA to learn eginespines from spine contours of training data, and reconstruct spine contours via LRA coefficients. To suppress inaccurate predictions and mask connectivity, we integrate LRA coefficient regression with anchor box classification to obtain spinal segment contours, in which sparse and dense assignments are integrated to more precisely detect spinal segments.
Privacy-Preserving Scoliosis Generation. We explore the latent diffusion model (LDM) [21] for scoliosis image synthesis. Specifically, we train LDM on our private dataset Spinal2023, utilizing a diffusion process to map raw images from pixel to latent space. During inference, the trained LDM with text-conditioned prompts can generate X-ray images. To mitigate privacy leaks [22], we employ data augmentation during LDM training, promote data balance, and introduce a privacy review [23] to exclude potentially privacy-compromising samples based on structural and pixel similarities, coupled with manual verification.
Cost-Effective Semi-Supervised Annotation. Inspired from semi-supervised pseudo-labeling [24] and multi-stage interactive annotation [25], we propose a new data engine, which drives an iterative pipeline with automatic annotation and selection. It consists of four stages: pseudo-labeling, auto-annotation, manual-assisted annotation, and privacy review. At each iteration, spine contour detection network is first re-trained using current dataset, and its trained model is used for pseudo-labeling so as to further update the dataset. This iterative pipeline culminates in our open-source dataset named Spinal-AI2024, which is the largest scoliosis X-ray dataset.
The main contributions of this paper are threefold:
  • We propose a novel low-rank approximation parameterized spine representation for automatic Cobb angle measurement, in which LRA coefficient regression and anchor box classification with sparse and dense assignments are combined to detect irregular contours of spinal segments.
  • We propose a new data engine with semi-supervised labeling and privacy review in an iterative manner, which enables the generation of large-scale scoliosis X-ray images largely without privacy leaks. To our knowledge, our open-source Spinal-AI2024 is the largest scoliosis X-ray dataset.
  • Extensive experiments on AASCE2019, Spinal2023, and Spinal-AI2024 datasets show that our method outperforms state-of-the-art automatic Cobb angle measurement works for both maximum angle and three regional angles.
2 Related Work
2.1 Scoliosis Cobb Angle Measurement
Earlier methods [10, 11, 26, 27] adopt traditional machine learning to approximate the relationship between image features and Cobb angles. However, these methods suffer from performance and generalization issues due to loss of high-resolution details through downsampling [17] or reliance on hand-crafted features. Recently, deep learning based methods [28, 29] have gained increasing attention.
Figure 2: The overall structure of our SG-LRA framework, which consists of spinal image generation, spinal contour detection, and data engine. Our private Spinal2023 dataset is used for training spinal image generation model, and also used for training initial model of spinal contour detection network. During each round of data engine, trained spinal contour detection network is adopted for pseudo-labeling, and selected data with annotations are backward used to fine-tune the spinal contour detection network. By employing the data engine to select and annotate generated images, we obtain the Spinal-AI2024 dataset.
Regression-based methods [12, 13, 27, 30] aim to learn a direct mapping between the spinal shape representation and the Cobb angle, without explicit geometric calculations. However, these approaches have limited performance for wedge-shaped or irregular vertebrae due to rectangular shape assumption.
Segmentation-based methods [14, 31, 32] usually rely on U-Net [33] or its modifications [34, 15], by performing pixel-level classification and then combining pixels as spinal regions. However, these methods are sensitive to image quality, and often require pre-processing to address noise, resolution, and contrast issues [35, 32].
Landmark-based methods [36, 16, 37, 17, 38] focus on localizing spinal landmarks, which often requires vast training data. In this paper, we overcome the data scarcity issue by predicting landmarks through LRA coefficient regression, leveraging our proposed data engine with automatic annotation and selection.
2.2 Medical Image Generation
Deep learning relies heavily on numerous labeled training data [39]. In the medical image field, there is often a scarcity of annotated data due to cumbersome, time-consuming, and expensive image acquisition and annotation [40]. To alleviate this limitation, previous works use data perturbation [41] or data synthesis [42].
Generative models like generative adversarial networks (GANs) [43, 44] have demonstrated the capability to produce realistic, high-resolution images [44, 45, 46, 47]. Recent studies [22] indicate that diffusion models [21] possess higher generation capacity than GANs. Packhäuser et al. [23] utilized LDM to generate chest X-ray images, and proposed a sampling strategy that safeguards the privacy of sensitive biometric information. However, sharing sensitive data under stringent privacy regulations remains a critical challenge in medical research [48, 49]. In this paper, we also leverage LDM to generate scoliosis X-ray images. Additionally, we propose a privacy-auditing module combined with pseudo-label training, utilizing a comprehensive privacy similarity mechanism to filter out images that may expose patient privacy.
2.3 Semi-Supervised Labeling
Semi-supervised labeling methods can be categorized into consistency regularization and pseudo-labeling approaches. Consistency regularization methods [50, 51] rely on modality-specific augmentation for regularization. Pseudo-labeling methods [52] predict labels for unlabeled data and train the model in a supervised manner.
Shi et al. [53] adopted local neighborhood density-based confidence scores to facilitate pseudo-labeling. Inspired by noise correction [54], Wang et al. [55] updated pseudo-labels by an optimization framework. Xie et al. [56] demonstrated the effectiveness of self-training in supervised classification tasks. However, these methods struggle with flawed pseudo-labels. Recently, Rizve et al. [24] selected high-quality pseudo-labels from noisy samples based on uncertainty. Kirillov et al. [25] built a data engine to augment data through manual-assisted, semi-automatic, and fully automatic phases. In this paper, we employ a similar semi-automatic annotation pipeline to expand datasets, in which pseudo-label selection and manual assistance are integrated for quality improvement.
3 Methodology
The architecture of our framework is illustrated in Fig. 2. Given our private Spinal2023 dataset, we first train Stable Diffusion [21] to generate vast spinal X-ray images. Specifically, the generation starts from the original pixel space, undergoing diffusion. The textual condition and pixel features are then fed into the latent space, where cross-attention modules produce the final output.
Also, we use Spinal2023 to train an initial model of spinal contour detection network. This network adopts ResNet50 [57] as backbone, and uses feature pyramid network (FPN) [58] to extract multi-scale features. Then, it uses regression and classification branches to predict LRA coefficients and positive anchor boxes, respectively. In both branches, sparse and dense assignments are combined to improve the detection performance. Finally, we select LRA coefficients corresponding to positive anchor boxes to reconstruct spinal segment contours.
The data engine integrates a loop with a screening mechanism. The screening process encompasses pixel-level privacy checks, sample-level legality reviews, and segment-level quality assessments. At each iteration, spine contour detection network is fine-tuned using current data, and its trained model is further used for pseudo-labeling so as to update the data and annotations.
3.1 Spinal Contour Detection via LRA
3.1.1 Parameterized Spine Representation
To address inaccuracy of rectangular box based spine representation in prior works, we utilize contour box composed by spinal landmarks to represent spinal segments. Given that spinal segments generally exhibit similar structural properties and notable correlations among their contours, we develop a new parameterized spinal contour representation method based on such correlations from training data.
In particular, the boundary of a spinal segment is composed of $N$ vertices, i.e. landmarks, which can be represented as a vector $\mathbf{a}=[x_{1},y_{1},\cdots,x_{N},y_{N}]\in\mathbb{R}^{2N\times 1}$ containing x and y coordinates. Then, we construct a spinal contour matrix $\mathbf{A}=[\mathbf{a}_{1},\mathbf{a}_{2},\cdots,\mathbf{a}_{L}]\in\mathbb{R}^{2N\times L}$ from training data containing $L$ spinal segment instances. Subsequently, we employ singular value decomposition (SVD) to extract the underlying structural relationship among $\mathbf{A}$:
$\mathbf{A=U\Sigma V}^{T}=\sigma_{1}\mathbf{u}_{1}\mathbf{v}_{1}^{T}+\cdots+\sigma_{r}\mathbf{u}_{r}\mathbf{v}_{r}^{T},$ (1)
where $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{2N}]\in\mathbb{R}^{2N\times 2N}$and $\mathbf{V}=[\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{L}]\in\mathbb{R}^{L\times L}$ are orthogonal matrices. $\Sigma\in\mathbb{R}^{2N\times L}$ is a diagonal matrix composed of singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}>0$, in which $r$ is the rank of $\mathbf{A}$.
Each singular vector represents an orthogonal direction in the eigen-space. By retaining the first $M$ singular values and their corresponding vectors while truncating the remaining low-rank components in Eq. (1), we can approximate $\mathbf{A}$ as
$\mathbf{A}_{M}=\sigma_{1}\mathbf{u}_{1}\mathbf{v}_{1}^{T}+\cdots+\sigma_{M}\mathbf{u}_{M}\mathbf{v}_{M}^{T}=\mathbf{U}_{M}\mathbf{\Sigma}_{M}\mathbf{V}_{M}^{T},$ (2)
where $\mathbf{A}_{M}$ is the best rank-M approximation of $\mathbf{A}$, since it minimizes the Frobenius norm $\|\mathbf{A}-\mathbf{A}_{M}\|_{F}$ [59].
Then, we define $\mathbf{C}_{M}=\mathbf{\Sigma}_{M}\mathbf{V}_{M}^{T}=[\mathbf{c}_{1},\mathbf{c}_{2},\cdots,\mathbf{c}_{L}]\in\mathbb{R}^{M\times L}$, such that $\mathbf{A}_{M}$ can be represented as
$\mathbf{A}_{M}=[\mathbf{U}_{M}\mathbf{c}_{1},\cdots,\mathbf{U}_{M}\mathbf{c}_{L}]=[\tilde{\mathbf{a}}_{1},\cdots,\tilde{\mathbf{a}}_{L}],$ (3)
where $\tilde{\mathbf{a}}_{i}=\mathbf{U}_{M}\mathbf{c}_{i}=[\mathbf{u}_{1},\mathbf{u}_{2},\cdots,\mathbf{u}_{M}]\mathbf{c}_{i}$ is the approximation of $\mathbf{a}_{i}$. We refer to $\mathbf{u}_{1},\ldots,\mathbf{u}_{M}$ as eigen-spines, since they are eigen-vectors of the matrix $\mathbf{AA^{T}}$ and can be viewed as principal spinal segment shapes.
For any $2N$-dimensional spinal segment contour $\mathbf{a}$, we can project it onto the eigen-spine space:
$\mathbf{c}=\mathbf{U}_{M}^{T}\mathbf{a},$ (4)
where $\mathbf{c}\in\mathbb{R}^{M\times 1}$ is the LRA coefficient vector.
3.1.2 Spinal Contour Detection Network
As shown in Fig. 2, our spinal contour detection network contains regression and classification branch, in which the overall loss is defined as
$\mathcal{L}=\lambda_{reg}\mathcal{L}_{reg}+\lambda_{cls}\mathcal{L}_{cls},$ (5)
where $\mathcal{L}_{reg}$ and $\mathcal{L}_{cls}$ are the losses for the regression and classification branches, respectively, and $\lambda_{reg}$ and $\lambda_{cls}$ control their importances. $\mathcal{L}_{cls}$ is composed by the spine region loss $\mathcal{L}_{sr}$ in dense assignment and the sparse sampling region loss $\mathcal{L}_{ssr}$ in sparse assignment:
$\mathcal{L}_{cls}=\mathcal{L}_{sr}+\mathcal{L}_{ssr},$ (6)
in which $\mathcal{L}_{sr}$ and $\mathcal{L}_{ssr}$ use the form of cross-entropy loss and focal loss, respectively. $\mathcal{L}_{reg}$ is defined as
$\mathcal{L}_{reg}=\sum_{i}\mathds{1}[i\in ps]l_{1}(\tilde{\mathbf{a}}_{i},\mathbf{a}_{i}),$ (7)
where $l_{1}$ denotes smooth-L1 loss, $ps$ denotes the positive sample region obtained by incorporating sparse and dense assignments, and $\mathds{1}$ denotes indicator function, outputting 1 when the condition is valid (the $i$-th point is within $ps$ in Eq. (7)) and outputting 0 otherwise. After combining the outputs of regression and classification branches, we can obtain a LRA coefficient matrix $\mathbf{C}_{K}\in\mathbb{R}^{M\times K}$. The predicted spinal segment locations are calculated as
$\mathbf{A}_{K}=\mathbf{U}_{M}\mathbf{C}_{K},$ (8)
where $K$ denotes the number of detected spinal segments.
Algorithm 1 The overall procedure of our data engine.
0:  Labeled dataset $D_{spinal2023}$, unlabeled dataset $D_{selected}$ generated by Stable Diffusion, sample selection index set $V_{selected}$ initialized as empty set, and spinal contour detection network with parameters $\Theta$.
0:  Labeled dataset $D_{SpinalAI2024}$.
1:  Train spinal contour detection network on $D_{spinal2023}$ and obtain initial parameters $\Theta^{0}$. Set $i\leftarrow 1$.
2:  while $V_{selected}$ is not converged do
3:     Pseudo-label samples in $D_{selected}$ using $\Theta^{i-1}$ with $\tau_{c}$-based positive spinal segment instance selection.
4:     Update $V_{selected}$ by auto-annotation, manual-assisted annotation, and privacy review.
5:     $D_{SpinalAI2024}\leftarrow\{s^{j}|s^{j}\in D_{selected},v^{j}=1\}$.
6:     Fine-tune network from $\Theta^{i-1}$ using samples from $D_{spinal2023}\cup D_{SpinalAI2024}$, and obtain new parameters $\Theta^{i}$. Set $i\leftarrow i+1$.
7:  end while
8:  Return $D_{SpinalAI2024}$.
3.2 Data Engine with Dataset Generation
Our data engine is composed of iteratively four stages: pseudo-labeling, auto-annotation, manual-assisted annotation, and privacy review. Before starting the data engine, we use Stable Diffusion [21] to generate an initial set of spinal X-ray images, as shown in Algorithm 1.
3.2.1 Pseudo-Labeling
At the $i$-th iteration, to enable pseudo-labeling, we first train our spinal contour detection network on currently available labeled data composed of $D_{spinal2023}$ and $D_{SpinalAI2024}$, in which the trained model parameters are denoted as $\Theta^{i-1}$. Then, we use $\Theta^{i-1}$ to label samples in $D_{selected}$. For the $j$-th sample $s^{j}$, the network predicts the locations of $K$ spinal segments $[\mathbf{a}_{1}^{j},\mathbf{a}_{2}^{j},\cdots,\mathbf{a}_{K}^{j}]$. To filter inaccurate predictions, we use a confidence threshold $\tau_{c}$ to select positive spinal segment instances for $s^{j}$:
$\mathcal{A}^{j}=\{\mathbf{a}_{k}^{j}|p^{j}_{k}\geq\tau_{c}\},$ (9)
where $p^{j}_{k}$ denotes the probability/confidence of the $k$-th spinal segment output by spinal contour detection network.
Table 1: The statistical characteristics of different datasets.
Dataset AASCE2019 (Public) Spinal2023 (Private) Spinal-AI2024 (Generated)
train test train test train test
Number of Images 481 128 700 204 16000 4000
Height of Images (Pixel) [973-3755] [1159-2880] [184-833] [227-802] [512-512] [512-512]
Width of Images (Pixel) [355-1427] [427-1386] [135-457] [147-415] [512-512] [512-512]
Cobb Angle (°) [5.96-90.0] [8.21-80.99] [3.22-74.26] [5.16-67.18] [0.0-90.0] [0.0-90.0]
Scoliosis Severity Num Avg Num Avg Num Avg Num Avg Num Avg Num Avg
Not ($<$ 10°) 17 8.35 1 8.21 91 7.82 20 8.24 1123 8.27 278 7.86
Mild (10° - 30°) 156 21.26 37 19.88 446 17.36 130 16.69 12268 18.39 2994 18.41
Moderate (30° - 45°) 156 37.53 47 38.74 76 37.87 31 37.56 1508 35.6 369 35.76
Severe ($>$ 45°) 152 60.89 43 56.98 87 55.87 23 56.1 1101 60.41 359 59.05
3.2.2 Auto-Annotation
We perform segment-level and sample-level automatic selection. We remove spinal segment instances with area smaller than 200 pixel2, remove instances with illegal landmark coordinates like negative values or positions beyond image boundaries, and remove instances whose landmark coordinates cannot form a valid contour. At the sample-level, we discard samples with fewer than 10 instances, in which the average number of spinal segments per image is 17 in real datasets. We also discard samples with spinal center distances larger than three times of the average distances. This sample-level selection can filter out samples with poor generation qualities. The index $v^{j}$ of sample $s^{j}$ is set to 1 in $V_{selected}$ if it is selected, and is set to 0 otherwise.
3.2.3 Manual-Assisted Annotation
This stage first adds annotations for samples with insufficient instances, and then manually corrects instances with inaccurate localization. We also remove non-realistic, spinal-fracture, and unclear samples.
3.2.4 Privacy Review
Finally, we remove generated samples with excessively high comprehensive similarity (CS) to real samples in $D_{spinal2023}$. The values of $V_{selected}$ are assigned to determine whether to discard certain samples:
$v^{j}=\mathds{1}[\max_{r=1}^{N_{r}}\{CS(s^{j},t^{r})\}\leq\tau_{CS}],$ (10)
where $CS(\cdot)$ denotes CS function, $t^{r}$ denotes the $r$-th sample in $D_{spinal2023}$, $N_{r}$ is the number of samples in $D_{spinal2023}$, and $\tau_{CS}$ is a threshold. We compare $s^{j}$ with all real samples so that the privacy of each real sample is preserved via $\tau_{CS}$-based selection. To assess the degree of patient privacy leakage in spinal image generation, we introduce the definition of memorization [22]: a sample $s^{j}$ is considered $N_{m}$-completely memorized if it can be extracted from the model and there are at most $N_{m}$ training samples $t^{r}$ in the dataset such that $CS(s^{j},t^{r})\leq\tau_{CS}$. Here, $N_{m}$ denotes the number of nearly identical samples to $s^{j}$, which can be close to zero by setting a strict $\tau_{CS}$ value.
To measure the similarity in terms of luminance, contrast, and structural information between two images, we use structural similarity index measure (SSIM):
$SSIM(s^{j},t^{r})\!=\!\frac{(2\mu_{s^{j}}\mu_{t^{r}}+c_{1})(2\sigma_{s^{j}t^{r}}+c_{2})}{(\mu_{s^{j}}^{2}+\mu_{t^{r}}^{2}+c_{1})(\sigma_{s^{j}}^{2}+\sigma_{t^{r}}^{2}+c_{2})},$ (11)
where $\mu_{s^{j}}$ and $\sigma_{s^{j}}^{2}$ are the mean pixel value and the variance of $s^{j}$, respectively, $\sigma_{s^{j}t^{r}}$ is the covariance of $s^{j}$ and $t^{r}$, and $c_{1}$ and $c_{2}$ are constants to stabilize division by denominators. Besides, we formulate the pixel similarity (PS) between gray-scale transformed $s^{j}$ and $t^{r}$ with sizes $H\times W$:
$PS(s^{j},t^{r})\!=\!\frac{1}{HWI}\sum_{a=1}^{H}\sum_{b=1}^{W}|GS(s^{j})_{ab}-GS(t^{r})_{ab}|,$ (12)
where $GS(\cdot)$ denotes gray-scale transformation, and $I$ denotes the number of gray scales (generally to be 255).
By combing Eqs. (11) and (12) using trade-off parameters $\lambda_{ss}$ and $\lambda_{ps}$, we obtain the definition of CS:
$CS(s^{j},t^{r})=\lambda_{ss}SSIM(s^{j},t^{r})+\lambda_{ps}PS(s^{j},t^{r}).$ (13)
After the above selection, we can obtain currently labeled generated dataset $D_{SpinalAI2024}$ using the selection index set $V_{selected}$. We then merge $D_{SpinalAI2024}$ with the existing $D_{spinal2023}$ to fine-tune the spinal contour detection network. We continue to generate pseudo-labels at the next iteration, and repeat this process until $V_{selected}$ converges.
4 Experiments
4.1 Datasets and Settings
4.1.1 Datasets
We evaluate Cobb angle measurement on public AASCE2019 [11, 60] benchmark, and our private Spinal2023 and open-source Spinal-AI2024 datasets. The dataset details are summarized in Table 1.
  • AASCE2019 contains 609 frontal spine X-ray images, in which each image is annotated by clinicians with 17 spinal segments and 68 landmarks, as well as PT, MT, and TL/L Cobb angles. This dataset exhibits a wide range of image intensities and Cobb angle variations. Following the official setting [60], we use 481 images for training and 128 images for testing.
  • Spinal2023 is our private clinical dataset, which consists of 904 spine X-ray images. It is randomly split into a training set of 700 images and a test set of 204 images. A public spinal landmark detection tool [17] is utilized to annotate 68 landmarks of 17 spinal segments in each image, which are further manually checked and corrected by clinical experts. The PT, MT, and TL/L Cobb angles are calculated based on spinal landmarks. Most images have mild or moderate scoliosis, and exhibit similar image intensities.
  • SpinalAI2024 is our generated dataset with totally 20,000 spine X-ray images, in which 16,000 and 4,000 images are used for training and testing, respectively. This dataset has a broad distribution of Cobb angles with a certain range of image intensities. The annotations of spinal landmarks as well as three Cobb angles for each image is obtained by our proposed data engine.
4.1.2 Implementation Details
Our SG-LRA is implemented via PyTorch. Similar to previous work [20], we set the eigen-spine space dimension $M$ and the spinal segment vertex number $N$ to 16 and $14$, respectively. The loss weights $\lambda_{reg}$ and $\lambda_{cls}$ in Eq. (5) are set to 0.1 and 1, respectively. We adopt stochastic gradient descent (SGD) optimizer and polynomial learning rate scheduler, with a weight decay of 0.0005, a momentum of 0.9, and an initial learning rate of 0.001. During pseudo-labeling, our spinal contour detection network is first pre-trained for 100 epochs on Spinal2023, and then are iteratively fine-tuned on training data composed of Spinal2023 images and selected generated images. The confidence threshold $\tau_{c}$ is set to 0.3. For privacy review, the weights $\lambda_{ss}$ and $\lambda_{ps}$ in Eq. (13) are set as 0.2 and 0.8, and the CS threshold $\tau_{CS}$ is set as 0.6.
4.1.3 Evaluation Metrics
Following the AASCE2019 [60] benchmark, we report symmetric mean absolute percentage error (SMAPE):
$\text{SMAPE}=\frac{100\%}{N}\times\sum_{i=1}^{N}\left(\frac{|Cobb_{pr}-Cobb_{gt}|}{|Cobb_{pr}|+|Cobb_{gt}|}\right),$ (14)
where $Cobb_{gt}$ and $Cobb_{pr}$ denote the ground-truth and predicted values, respectively. It is used to evaluate the accuracy of the maximum Cobb angle, as well as each Cobb angle of PT, MT, and TL/L. Besides, we use object detection metrics average precision (AP) and average recall (AR), as well as angle difference metric Euclidean distance (ED) to assess the effect of iterative label refinement of spinal contours.
Table 2: Comparison with state-of-the-art methods on AASCE2019 in terms of maximum Cobb angle evaluation. These methods are divided into traditional, regression-based, segmentation-based, and landmark-based methods, respectively. Our SG-LRA can be regarded as combining regression-based and landmark-based methods.
       Method        SMAPE (%)
       S2VR [10]        37.08
       BoostNet [11]        23.44
       MVC-Net [27]        35.85
       Faster-RCNN [12]        25.7
       MVE-Net [30]        18.95
       ResNet [13]        10.81
       Seg4Reg [14]        21.71
       Residual U-Net [34]        16.48
       TOLACT [32]        10.76
       Augmented U-Net [15]        9.2
       CenterNet [37]        13.46
       LDNet [17]        10.81
       HrNet [36]        10.03
       Hourglass [16]        9.78
       Linformer [38]        7.91
       SG-LRA        5.05
4.2 Comparison with State-of-the-Art Methods
4.2.1 Maximum Cobb Angle Evaluation
Table 2 shows comparison results on the public AASCE2019 dataset. It can be seen that our SG-LRA maintains top-tier performance on AASCE2019. Compared to regression-based methods like Faster-RCNN, SG-LRA performs better, due to our LRA parameterized spine representation for precise Cobb angle calculation. Compared to segmentation-based methods, SG-LRA with sparse and dense assignments can more accurately detect spinal segment contours, better addressing issues of spinal mask connectivity/fragmentation. Compared to landmark-based methods, our LRA coefficient regression combined with the anchor box classification branch can more accurately localize corner points, giving stronger robustness for non-rectangular spinal segments.
Figure 3: Visualization of different spine representations on example images from AASCE2019 and Spinal2023.
Besides, we visualize different spine representations in Fig. 3. It can be seen that our LRA parameterized spine representation is more fine-grained than landmark-based and rectangular-box-based representations, enabling more precise Cobb angle calculation.
Table 3: SMAPE (%) results in terms of maximum Cobb angle and three regional Cobb angles on different datasets. The results on AASCE2019 are reported by literatures, and those on Spinal2023 and Spinal-AI2024 are implemented using their code.
Method AASCE2019 (Public) Spinal2023 (Private) Spinal-AI2024 (Generated)
Max PT MT TL/L Max PT MT TL/L Max PT MT TL/L
LDNet 10.81 6.26 18.04 23.42 17.22 23.88 24.12 32.62 15.23 32.68 21.1 27.57
Hourglass 9.78 11.13 15.22 21.03 16.05 21.63 26.67 24.5 10.25 26.8 18.6 20.02
Residual U-Net 16.48 9.71 25.97 33.01 24.3 16.6 31.2 35.4 13.3 10.79 19.93 38.18
Linformer 7.91 4.86 14.65 20.49 19.1 17.25 21.17 29.17 11.59 7.83 18.82 24.62
SG-LRA 5.05 12.48 6.27 14.04 7.13 23.3 9.49 17.25 5.34 22.91 8.64 16.69
4.2.2 PT, MT, and TL/L Cobb Angle Evaluation
We conduct a more detailed comparison by evaluating Cobb angles of three regional spinal curves in Table 3. Since the construction of Spinal-AI2024 involves our spinal contour detection network, the reported results of SG-LRA on Spinal-AI2024 are obtained by re-training spinal contour detection network on Spinal-AI2024. It can be observed that our SG-LRA significantly outperforms previous works in most cases.
In particular, on all the three datasets, SG-LRA performs the best on MT and TL/L Cobb angles, as well as maximum Cobb angle. This is mainly thanks to the combination of LRA coefficient regression and anchor box classification. Note that the performance for PT cure is not as good, possibly because SG-LRA focuses more on localizing irregular spinal segments rather than limiting the number of predicted spinal segments. As a result, it may involve the cervical spine in the calculation of PT Cobb angle, leading to worse results. Although previous methods fix the predicted whole scoliosis to 17 spinal segments, which often obtains more accurate PT predictions, complex pre-processing of data and annotations based on clinical knowledge are required.
It can be seen that segmentation-based method Residual U-Net achieves the best PT Cobb angle estimation result on Spinal2023, and landmark-based method Linformer achieves the best PT Cobb angle estimation result on Spinal-AI2024. We notice that Spinal2023 has limited images while Spinal-AI2024 has a larger scale. In this case, landmark-based methods suffer from insufficient training data on Spinal2023 but can be trained well on Spinal-AI2024, while segmentation-based methods have larger advantages on Spinal2023 with smaller scale. In contrast, our SG-LRA consistently works well across small and large datasets.
Table 4: Maximum Cobb angle estimation results of different variants of our spinal contour detection network on Spinal2023. $C_{sparse}$, $C_{dense}$, $R_{sparse}$, and $R_{dense}$ denotes sparse assignment and dense assignment in the classification and regression branches, respectively.
$C_{sparse}$ $C_{dense}$ $R_{sparse}$ $R_{dense}$ SMAPE(%)
61.03
53.97
25.02
17.25
7.13
4.3 Ablation study
In this section, we evaluate the key modules in our framework: spinal contour detection network and data engine.
4.3.1 Spinal Contour Detection Network
Our spinal contour detection network consists of LRA coefficient regression branch and anchor box classification branch, with each branch containing sparse and dense assignments. Table 4 presents the results of different variants on Spinal2023.
Regression and Classification Branches. We can see that using classification or regression branches alone both perform poor, in which the regression is more important. When combing both regression and classification, our full spinal contour detection network achieves the smallest SMAPE.
Sparse and Dense Assignments. In the case of combining both branches, only utilizing sparse or dense assignments still cannot achieve the desired Cobb angle measurement effect. Due to more fine-grained prediction, the use of dense assignment has a higher accuracy than sparse assignment.
Figure 4: Performance of our spinal contour detection network under different confidence thresholds on generated dataset obtained at the first data engine iteration. Note that when $\tau_{c}$ is set optimally at 0.3, ED and SMAPE results on the final Spinal-AI2024 dataset will be increased from 20.73 and 26.13% to 6.23 and 5.34%, respectively.
4.3.2 Data Engine
During our framework, the X-ray images generated by Stable Diffusion [21] have a certain degree of noises, in which some severe noises do not conform to reality. This problem is widespread in medical image generation, due to the distribution characteristics of training data itself, including the lack of rare cases like Cobb angles greater than 45 degrees. Our data engine is proposed to solve this issue, and is validated in this section.
Different Confidence Thresholds. To investigate the effects of different confidence thresholds, we train spinal contour detection network on $D_{SpinalAI2024}$ obtained at the first data engine iteration using different $\tau_{c}$ with a interval of 0.1, as illustrated in Fig. 4. The results show that setting $\tau_{c}$ to 0.3 results in the highest pseudo-label qualities. When $\tau_{c}$ is lower, the selected spinal segment instances may contain more noises, which complicates the further denoising process. When $\tau_{c}$ is higher, important positive instances may be missing, causing large errors in calculating Cobb angles.
Table 5: AP and AR results on six subsets of Spinal-AI2024 before and after denoising at the first data engine iteration.
Subset AP AR
Before After ($\uparrow$) Before After ($\uparrow$)
1 60.5 79.8 (19.3) 15.9 20.8 (4.9)
2 80.1 89.9 (9.8) 20.9 26.3 (5.4)
3 89.6 94.1 (4.5) 26.3 32.9 (6.6)
4 95.2 96.5 (1.3) 33.0 39.8 (6.8)
5 94.1 95.5 (1.4) 32.9 38.8 (5.9)
6 95.6 96.6 (1.0) 39.0 45.9 (6.9)
Table 6: Results using different selections on Spinal-AI2024 from the first to the six iterations. “No”, “Indep.”, and “Cumul.” denote no filtering, and updating $V_{selected}$ in independent and cumulative ways across iterations, respectively.
Selection 1 2 3 4 5 6
No 84.0 87.9 95.3 96.4 96.6 94.8
AP Indep. 84.3 88.1 95.5 96.5 96.6 96.7
Cumul. 84.3 89.2 95.6 96.6 97.8 96.9
No 26.0 32.1 38.8 45.5 56.1 72.8
AR Indep. 26.1 31.8 38.8 45.6 56.2 74.3
Cumul. 26.1 32.0 38.8 45.6 56.2 74.4
Denoising in Auto- and Manual-Assisted Annotation. Table 5 shows the results before and after denoising, in which AP and AR are calculated between current labels and the labels in the final Spinal-AI2024. It is indicated that denoising significantly improves the qualities of pseudo labels.
Different Sample Selection Strategies. We conduct three experiments using different sample selection strategies, as shown in Table 6. If a sample is filtered and will not be selected in the next iterations, i.e. cumulative way, the overall highest quality improvement can be achieved. If $V_{selected}$ is updated independently across iterations, the quality improvement becomes worse but is still better than no filtering. This demonstrates the cumulative effectiveness of combined denoising and privacy review processes.
We also present the filtered spinal segment instances and unreasonable samples during sample selection in Fig. 5. It can be seen inaccurate pseudo-labels and poor generated samples are discarded.
5 AI Responsibility and Ethical Standards
The aim of this research is to enhance the accuracy and safety of automatic Cobb angle measurement for scoliosis diagnosis. Careful data scrutiny and adherence to ethical standards are crucial when introducing this technology. A thorough privacy review of the AI-generated data is necessary to protect patient information and rights. Upholding the highest ethical principles is a fundamental responsibility for researchers in this domain.
Figure 5: Illustration of discarded spinal segment instances as well as unreasonable samples during sample selection. (a) Inaccurate spinal segment instances due to low threshold; (b) Spinal segment instances with area smaller than 200 pixel2; (c) Samples with fewer than 10 instances; (d) Image distortion; (e) Spinal fracture; (f) Samples with excessively high comprehensive similarity (CS).
[…]
New Image Top1-Image Top1-SSIM $\cdots$ Top1-Image Top1-PS $\cdots$ Top1-CS Top2-CS Top3-CS ACS
000001.jpg 000196.png 38.24 $\cdots$ 000290.png 37.78 $\cdots$ 47.61 47.15 46.77 47.18
001257.jpg 000527.png 52.76 $\cdots$ 000427.png 88.58 $\cdots$ 59.92 58.18 58.01 58.71
020140.jpg 000009.png 44.2 $\cdots$ 000531.png 89.3 $\cdots$ 53.22 53.15 52.83 53.07
020000.jpg 000763.png 34.46 $\cdots$ 000493.png 82.25 $\cdots$ 44.02 43.91 43.88 43.93
5.1 Data Availability
The datasets used in this paper include the public AASCE2019 benchmark, our private Spinal2023 dataset, and our generated Spinal-AI2024 dataset. Spinal-AI2024 has been released at https://github.com/Ernestchenchen/Spinal-AI2024. Although Spinal2023 can not be made publicly available due to patient privacy, this paper has provided many details regarding the data source, data processing, and data distribution. Besides, Spinal2023 is granted under a strict data usage agreement and is approved by the institutional Medical Research Ethics Committee. Spinal-AI2024 conducts privacy assessments to prevent patient re-identification.
5.2 Data Ethics
5.2.1 Public AASCE2019
The use of this benchmark dataset allows objective evaluation and comparison of our automatic Cobb angle measurement techniques. Importantly, the AASCE dataset has undergone appropriate anonymization and de-identification to protect patient privacy, in alignment with relevant ethical guidelines and regulations.
5.2.2 Private Spinal2023
In addition to the public AASCE dataset, this research also utilized the private Spinal2023 dataset, which contains X-ray images and Cobb angle annotations for a larger scoliosis patient cohort. Access to Spinal2023 was granted under a strict data usage agreement with comprehensive privacy and protection protocols, reviewed and approved by the institutional Medical Research Ethics Committee.
Figure 6: Illustration of highest similar images.
5.2.3 Generated Spinal-AI2024
This dataset was created by training generative AI models on Spinal2023 to synthesize additional X-ray images and Cobb angle annotations. The research team carefully reviewed and validated the Spinal-AI2024 dataset to ensure it preserves the original data characteristics.
The inclusion of the Spinal-AI2024 dataset expands the training and evaluation data, which can enhance the robustness and generalizability of the automatic Cobb angle measurement algorithms. However, the use of AI-generated data requires additional ethical considerations. We have implemented rigorous data provenance tracking, algorithmic transparency, and human oversight to verify the Spinal-AI2024 dataset’s integrity and fidelity. Comprehensive privacy assessments have also been conducted to prevent potential patient re-identification.
For each Spinal-AI2024 image, a top-3 comprehensive similarity review is presented in Table 7, in which the most similar images are shown in Fig. 6. The results show that our generated dataset has distorted the original annotations and produced spinal curvature variations, which can prevent excessive patient information leakage.
6 Conclusion
We have proposed a novel self-generating framework utilizing LRA parameterized spine representation for automatic Cobb angle measurement. LRA coefficient regression and anchor box classification with sparse and dense assignments are combined to detect irregular contours of spinal segments. Besides, we have proposed a data engine with semi-supervised labeling and privacy review, which culminates in the open-source largest scoliosis X-ray dataset Spinal-AI2024. Extensive experiments have demonstrated the effectiveness of our method, achieving state-of-the-art performance on public, private, and generated datasets for both maximum Cobb angle and three regional Cobb angles.